Kroger's 2012 financial statements show average shareholders' equity of $4,632 million, net income of $602 million, interest expense of $435 million, and average total assets of $23,491 million.

How much is Kroger's return on assets for the year? Assume that the statutory rate is 35%.

A.

10.7%

B.

0.6%

C.

3.8%

D.

9.4%

E.

There is not enough information to calculate the ratio
Answer #1

Return on assets = (Net income+Interest net of tax)/Average total assets

Interest expense (net of tax) = 435*65% = 282.75

Return on assets = (602+282.75)/23491 = 3.8%

So answer is c) 3.8%